OpenStudy (anonymous):

HELP PLEASE PRECAL I NEED A GOOD GPA OK.. WHICH OF THE FOLLOWING EQUATIONS WILL PRODUCE THE GRAPH SHOWN BELOW?

OpenStudy (anonymous):

yes i am taking a test but i'm failing the class and it's important to me

OpenStudy (anonymous):

a) \[x ^{2}-\frac{ y ^{2} }{ 4}=1\] b) \[\frac{ y ^{2} }{ 9} -\frac{ x ^{2}\ }{ 4 }\]

OpenStudy (anonymous):

c) \[y ^{2} - \frac{ x^2 }{ 9}\]

OpenStudy (anonymous):

d) \[\frac{ y^2 }{ 2 } - \frac{ x^2 }{ 4 } =1\]